tickets for a school playsell for $8for floor seatsand $6 for balcony seats.For the performance 72 tickets are sold,bringing in $516.how many of each tickets are sold?
Answer by checkley77(12844)

F+B=72 or F=72-B
8F+6B=516
8(72-B)+6B=516
576-8B+6B=516
-2B=516-576
-2B=-60
B=-60/-2
B=30 BALCONY TICKETS WERE SOLD.
F=72-30
F=42 FLOOR SEATS WERE SOLD.
PROOF:
8*42+6*30=516
336+180=516
516=516